OpenStudy (anonymous):

Astronauts on our moon must function with an acceleration due to gravity of 0.170g .If an astronaut can throw a certain wrench 15.0m vertically upward on earth, how high could he throw it on our moon if he gives it the same starting speed in both places?

OpenStudy (carmuz):

PHYSICAL APPROCH: You know that \[{V_f}^2 = {V_0}^2 - 2gh\] The wrench is going upwards and when it reach the top of its path, Vf = 0. Therefore, you will have: \[{V_0}^2 = 2gh    \rightarrow    V_0 = \sqrt{2gh}\] Vo is the same in both places, so you will have: \[{V_0}^2 = 2g_{earth}h_{earth}\] and \[{V_0}^2 = 2g_{moon}h_{moon}\] then \[g_{earth}h_{earth} = g_{moon}h_{moon}\]but \[g_{moon}= 0.17g_{earth}\] then \[g_{earth}h_{earth} = 0.17g_{earth}h_{moon}\]\[h_{earth} = 0.17h_{moon}\] Finally: \[15 = 0.17h_{moon}\] \[h_{moon} = 88.2 m\] Mathematical Approach: Height is inversely proportional to gravity. \[\frac{ g_{earth} } { g_{moon} } = \frac{ h_{moon} }{ h_{earth} }\]You will get the same height; h = 88.2m.

OpenStudy (anonymous):

thank u so much